# Distillation Improves Visual Place Recognition for Low Quality Images

### v2

@@ IV-A2 Testing Datasets
We tested $f_{s}$ after distillation using each loss combination by calculating its VPR recall rate on three VPR datasets organized using VPR Datasets Downloader, namely Mapillary SLS [29], Nordland [26], and Tokyo 24/7 [28]. For each dataset, we follow the designated test partition and its database and query image specifications. Additional context for each dataset is provided below:
  • Mapillary SLS is specifically curated for lifelong VPR featuring images from 30 cities.
  • Nordland is created with forward-facing recordings across four seasons from trains along a rail route. Our version of the dataset is sourced identically as [9].
  • Tokyo 24/7 depicts urban scenes at different times of day within Tokyo’s Shibuya ward.
Table II: Best Performance of VPR Methods after Distillation: For low quality images, the loss combination producing the highest recall rate is compared against the baseline of fine-tuning. For both distilled and fine-tuned weights, the change in VPR recall is represented as a delta relative to the performance of each method using pretrained weights. Within each dataset and method, greene text indicates the greatest improvement for every R@N, whereas any decrease relative to pretrained performance is marked as red. The recall rates using pretrained weights on unmodified $I^{h}$ images are provided as reference.
@@ IV-B VPR Methods
To validate the applicability of knowledge distillation for VPR, we carried out the aforementioned experiments on five VPR methods from various points in the VPR literature, namely NetVLAD [8], MixVPR [9], AnyLoc [10], DINOv2SALAD [17], and CricaVPR [19]. While the former two methods have CNN backbones, the remaining methods are based on DINOv2 [44]. The diverse architectures of these five methods allow us to explore our distillation approach’s applicability to diverse VPR methodologies. Furthermore, to study each loss component’s effects on $f_{s}$’s performance, we trained $f_{s}$ under each of the seven possible combinations of $L_{\theta 1}$, $L_{\theta 2}$, and $L_{\theta 3}$, using at least one term from (4). For each VPR method, we acquire pretrained model weights released with their public source codes to act as $f_{t}$ and initialize $f_{s}$ for faster training convergence.
As a baseline to be compared against distillation, we also fine-tune all methods except for AnyLoc using their unmodified training source code on the same $I^{l}$ and pretrained weights as distillation. Given that AnyLoc uses off-the-shelf DINOv2 backbone and does not learn its aggregator weights, we only apply distillation to this method.
@@ IV-C1 Training Configuration
We set the loss weighting coefficients in (4) to $\alpha=10^{5}$ and $\beta=10^{4}$. The learning rate is initialized at $10^{-5}$, modulated by a decay factor of $2\times 10^{-11}$, and further adjusted by an exponential rate of $0.99999$. Each training experiment (ie. one of seven loss combinations with one of five methods) consists of one complete iteration through the shrunken GSV-Cities dataset IV-A1 (ie. one epoch). For the triplet loss (3), we sampled $5$ negative images to form $v^{l}_{n}$ for each training image.
@@ IV-C2 VPR Performance Evaluation
To assess $f_{s}$’s VPR performance after distillation, we use the metric Recall at N (referred to as R@N) found in [8, 9, 16, 7, 17, 10, 19] to calculate $f_{s}$’s recall rate, which measures the percentage of successfully localized query images. Given a choice of $N$, a query to the VPR database is deemed successful if one of the top $N$ retrieved database images is within $d$ meters of the query’s true position for some predetermined threshold $d$. The percentage of successfully retrieved query images becomes $f_{s}$’s R@N score for that $N$. In our evaluation procedures, we selected $N$ values of $1$, $2$, $5$, and $10$, whereas $d$ remained fixed at 25 meters following the distance thresholds used by the VPR methods we experimented with.
@@ V Results and Discussions
For each loss combination, We use $f_{s}$ after distillation to generate $v^{l}$ for the selected testing datasets IV-A2. We then calculate recall rates for different versions of $f_{s}$ and analyze the efficacy of each configuration of our proposed distillation methodology.
@@ V-A Quantitative Results
Table II demonstrates the applicability of (4) and its components to a representative selection of VPR methods and datasets. Albeit with varying magnitudes, the deltas of recall rates under $v^{l}$ after distillation and those under $v^{l}$ from pretrained weights are generally positive. However, fine-tuning is inconsistent and mostly less effective than distillation. The Nordland recall rates for MixVPR and CricaVPR, two architecturally distinct methods, exemplifies distillation’s superior efficacy and applicability. There are only two instances where distillation partially failed to increase VPR recall rates under $v^{l}$ for some method and dataset, namely NetVLAD on Nordland and AnyLoc on Tokyo 24/7. As AnyLoc’s recall rates fluctuate greatly on Tokyo 24/7, we postulate the combination of the natures of this method and dataset to be an edge case. Since NetVLAD’s improvement on Nordland for all R@N is less than that on other datasets, we reserve this phenomenon for further analysis in V-B.
@@
V-B Qualitative results
Table III: Indoor Evaluation Using Pretrained Weights
@@ VI Additional Experiments
While our knowledge distillation approach already demonstrates notable efficacy for the majority of methods and datasets tested, we extend distillation to more datasets and other aforementioned routes of image quality reduction II-A.
@@ VI-A JPEG Compression’s Impacts on Indoor Datasets
While our previous VPR testing datasets cover both rural and urban environments and contain temporal changes and perceptual aliasing, they all capture outdoor environments. In table III, We further examine the impacts of JPEG compression on three indoor datasets of different difficult levels, namely TUM LSI [27], Gangnam Station (one of NAVER LABS’ large-scale localization datasets [30]), and NYC-Indoor-VPR [33]. While the performance of all methods with pretrained weights fluctuate greatly across datasets, JPEG compression’s impacts on recall rates are significantly smaller than previous results with outdoor datasets. We postulate that due to indoor environments generally lack large contiguous regions of uninformative features such as sky, the proportion of distinctive features degraded by JPEG compression is smaller within indoor images. Therefore, instead of repeating our previous experiments, we extend our distillation methodology to other possible forms of $I^{l}$.
@@ VI-B Other Modalities of Image Quality Reduction
Inferring from JPEG compression’s greater impacts on outdoor scenes, we use outdoor datasets from IV-A2 to yield $I^{l}$ with reduced resolution instead of JPEG compression. Secondly, we explore the impacts of video quantization on VPR recall, mirroring realistic practice of streaming image data through video. However, the image-based nature of existing VPR datasets prompts us to curate a custom video-based VPR dataset to produce $I^{l}$ with increased video quantization.
Amongst the five methods in table II, NetVLAD suffers the most from lowered image quality in general. Thus, we select NetVLAD as the candidate of verifying knowledge distillation’s benefits for VPR performance under resolution and quantization-based $I^{l}$.
@@ VI-D Resolution and Quantization Experimental Results
For outdoor datasets, we compare pretrained NetVLAD against trained $f_{s}$ with the highest recall rate in table IV. NetVLAD’s best-performing distillation loss remains MSE, but the addition of ICKD yields stronger performance on Tokyo 24/7. This corroborates the previous observation of ICKD’s adaptability, which may allow $f_{s}$ to better withstand changing $I^{l}$ modalities from $180$p during distillation to $240$p in testing.
Table IV: NetVLAD’s Performance on $240$p Images
@@ Acknowledgment
We would like to express our gratitude to Zezheng Li and Liyuan Geng for their valuable assistance with test data preprocessing. Their contributions helped facilitate the progress of this work.

### v3

@@ IV-A2 Testing Datasets
We evaluate $f_{s}$ on three standard VPR datasets: Mapillary SLS[21], Nordland[22], and Tokyo 24/7[20]
Table I: Best Performance of VPR Methods after Distillation: For low quality images, the loss combination producing the highest recall rate is compared against the baseline of fine-tuning. For both distilled and fine-tuned weights, the change in VPR recall is represented as a delta relative to the performance of each method using pretrained weights. Within each dataset and method, greene text indicates the greatest improvement for every R@N, whereas any decrease relative to pretrained performance is marked as red. The recall rates using pretrained weights on unmodified $I^{h}$ images are provided as reference.
@@ IV-B VPR Methods
To validate the broad applicability of our knowledge distillation framework, we experiment with five VPR methods representing different architectures:
  • NetVLAD [16] and MixVPR [6] (CNN-based),
  • AnyLoc [34], DINOv2SALAD [35], and CricaVPR [7] (DINOv2-based [36]).
For each method, we initialize both the teacher $f_{t}$ and student $f_{s}$ using pretrained weights released by each method’s authors for faster convergence. To investigate the effects of individual loss components, we train $f_{s}$ under all seven possible combinations of $L_{\theta 1}$, $L_{\theta 2}$, and $L_{\theta 3}$ (with at least one term used from Eq. (4)).
As a baseline, we also fine-tune each method (except AnyLoc, which uses a frozen DINOv2 backbone) on the same $I^{l}$ images using pretrained weights and the authors’ provided training procedure instead of our distillation losses. This baseline comparison highlights the improvement attributable solely to our distillation framework.
@@ IV-C1 Training Configuration
Our loss weighting coefficients in Eq. (4) are set to $\alpha=10^{5}$ and $\beta=10^{4}$. These values were chosen based on preliminary ablation studies aimed at balancing the scale of the MSE and triplet losses relative to the ICKD loss. The learning rate is initialized at $10^{-5}$ and adjusted using a decay factor of $2\times 10^{-11}$ along with an exponential decay rate of $0.99999$. Each training experiment (one of the seven loss combinations applied to one of the five VPR methods) comprises a single epoch over the reduced GSV-Cities dataset (Section IV-A1). For the triplet loss (Eq. (3)), we sample 5 negative images per training sample to form $v^{l}_{n}$.
@@ IV-C2 VPR Performance Evaluation
We assess the VPR performance of the distilled student network $f_{s}$ using the Recall at N (R@N) metric, following prior work [16, 6, 37, 10, 35, 34, 7]. R@N is defined as the percentage of queries for which at least one of the top $N$ retrieved database images is within a specified distance threshold $d$ of the query’s true location. Following the VPR methods we experimented with, we set $d$ to $25$ meters, and we evaluate R@N for $N$ values of 1, 2, 5, and 10.
@@ V Results and Discussions
We analyze the performance of the student network $f_{s}$ after distillation using each loss combination. We report quantitative improvements in VPR recall rates, discuss observed trends across different methods and datasets, and provide qualitative visualizations to interpret the behavior of the network after distillation.
@@ V-A Quantitative Results
Table I demonstrates the applicability of (4) and its components to a representative selection of VPR methods and datasets. Albeit with varying magnitudes, the deltas of recall rates under $v^{l}$ after distillation and those under $v^{l}$ from pretrained weights are generally positive. However, fine-tuning is inconsistent and mostly less effective than distillation. The Nordland recall rates for MixVPR and CricaVPR, two architecturally distinct methods, exemplifies distillation’s superior efficacy and applicability. There are only two instances where distillation partially failed to increase VPR recall rates under $v^{l}$ for some method and dataset, namely NetVLAD on Nordland and AnyLoc on Tokyo 24/7. As AnyLoc’s recall rates fluctuate greatly on Tokyo 24/7, we postulate the combination of the natures of this method and dataset to be an edge case. Since NetVLAD’s improvement on Nordland for all R@N is less than that on other datasets, we reserve this phenomenon for further analysis in V-B.
@@
Table II: Indoor Evaluation Using Pretrained Weights
@@ VI Additional Experiments
While our knowledge distillation approach already demonstrates notable efficacy for the majority of methods and datasets tested, we extend distillation to more datasets and other modes of image quality reduction discussed in II-A.
@@ VI-A JPEG Compression’s Impacts on Indoor Datasets
While our previous VPR testing datasets cover both rural and urban environments and contain temporal changes and perceptual aliasing, they all capture outdoor environments. In table II, We further examine the impacts of JPEG compression on three indoor datasets of different difficulty levels, namely TUM LSI [38], Gangnam Station (one of NAVER LABS’ large-scale localization datasets [39]), and NYC-Indoor-VPR [40]. While the performance of all methods with pretrained weights fluctuate greatly across datasets, JPEG compression’s impacts on recall rates are significantly smaller than previous results with outdoor datasets. We postulate that due to indoor environments generally lack large contiguous regions of uninformative features such as sky, the proportion of distinctive features degraded by JPEG compression is smaller within indoor images. Therefore, instead of repeating our previous experiments, we extend our distillation methodology to other possible forms of $I^{l}$.
@@ VI-B Other Modalities of Image Quality Reduction
As JPEG compression had greater impacts on outdoor scenes, we use outdoor datasets from IV-A2 to yield $I^{l}$ with reduced resolution instead of JPEG compression. Secondly, we explore the influence of video quantization on VPR recall, mirroring realistic practice of streaming image data through video. However, the image-based nature of existing VPR datasets prompts us to curate a custom video-based VPR dataset to produce $I^{l}$ with increased video quantization.
Among the five methods in table I, NetVLAD suffers the most from lowered image quality in general. Thus, we select NetVLAD as the candidate of verifying knowledge distillation’s benefits for VPR performance under resolution and quantization-based $I^{l}$.
@@ VI-D Resolution and Quantization Experimental Results
For outdoor datasets, we compare pretrained NetVLAD against trained $f_{s}$ with the highest recall rate in table III. NetVLAD’s best-performing distillation loss remains MSE, but adding ICKD yields stronger performance on Tokyo 24/7. This corroborates the previous observation of ICKD’s adaptability, which may allow $f_{s}$ to better withstand changing $I^{l}$ modalities from $180$p during distillation to $240$p in testing.
Table III: NetVLAD’s Performance on $240$p Images
@@ Acknowledgment
We would like to express our gratitude to Zezheng Li and Liyuan Geng for their valuable assistance with test data preprocessing, accelerating this work’s progress.